A $1,000 par bond with a 8% semi-annual coupon and a yield to maturity of 5.5% has 15 years to maturity. What is the price of the bond?

  • A. $1,250.94

  • B. $781.99

  • C. $1,253.12

  • D. $984.33
Answer #1

Price of a bond is mathematically represented as: 1 1-- (1 + i) PEC. - +- м. (1 + i) where P is price of bond, with periodiFor this question,

M = $1000, n = 15 * 2 = 30 semi-annual periods,i = 5.5%/2 = 2.75% (semi-annually), C = 8% * $1000/2 = $40 (semi-annually)

1- (1+0.0275)30 P = 40 * - 0.0275 1000 (1 + 0.0275) 30

P = $809.97 + $443.14

P = $1,253.12